# supplement

### v1

@@ 1 Introduction
In demography, the age distribution of deaths provides important insights into longevity and lifespan variability that cannot be grasped directly from either the central mortality rate or the survival function. For modelling the age distribution of deaths, Oeppen (2008) demonstrated that using compositional data analysis (CoDa) to forecast the age distribution of death counts does not lead to more pessimistic results than forecasting age-specific mortality. Thus, the age distribution of death counts is more suitable than the central mortality rates for computing period life expectancy and annuity values. In addition to providing an informative description of the mortality experience of a population, the age-at-death distribution yields readily available information on the “central longevity indicators” (e.g., mean, median and modal age at death, see Cheung et al. (2005) and Canudas-Romo (2010)) and lifespan variability (e.g., Robine, 2001; Vaupel et al., 2011; van Raalte & Caswell, 2013; van Raalte et al., 2014; Aburto & van Raalte, 2018).
@@
Using the Japanese age- and sex-specific life-table death counts from 1975 to 2022, we evaluate and compare one- to 20-step-ahead point and interval forecast accuracy between the CoDa and CDF transformations. To evaluate point forecast accuracy, we use the Kullback-Leibler divergence and the Jensen-Shannon divergence. To assess the interval forecast accuracy, we utilise the interval score of Gneiting & Raftery (2007) and Gneiting & Katzfuss (2014); see Section 4 for details. The improved forecast accuracy of life-table death counts is of great importance to demographers and actuaries for determining remaining age-specific life expectancies and to actuaries for pricing temporary annuities for various ages and maturities.
@@ 2 Age- and sex-specific life-table death counts in Japan
In Figure 1, we display Japanese age- and sex-specific life-table death counts from 1975 to 2022, obtained from the Japanese Mortality Database (2024). We use data from the period after the First and Second World Wars to obtain a more stable parameter estimate from the historical data. We study life-table death counts, where the life-table radix (i.e., a population experiencing 100,000 births annually) is fixed at 100,000 at age 0 while the remaining number of people alive is 0 in the last age group for each year. For the life-table death counts, there are 111 ages, and these are ages 0, 1, …, 109, 110+. Due to rounding, there are zero counts for age 110+ at some years. To overcome this problem, we work with the probability of dying (i.e., $q_{x}$) and the life-table radix to recalculate our estimated death counts (up to six decimal places). In doing so, we obtain more detailed death counts than the ones reported in the Japanese Mortality Database (2024).
@@
Figure 1: Rainbow plots of the age distribution of deaths from 1975 to 2022 in a single-year group. The life-table radix is 100,000 for each year. The life-table death counts in the oldest years are shown in red, while the most recent years are in violet. Curves are ordered chronologically according to the colours of the rainbow.
The spread of the distribution indicates lifespan variability. A decrease in variability over time can be observed directly and can be measured, for example, with the interquartile range of life-table ages at death or the Gini coefficient (for comprehensive reviews, see Wilmoth & Horiuchi, 1999; van Raalte & Caswell, 2013; Debón et al., 2017). For modelling the age distribution of deaths, a Gini coefficient of 0 indicates perfect inequality among ages in life-table death counts, and a value of 1 indicates perfect equality, implying that death occurs at the ultimate same age. From Figure 2, the female series has a higher Gini coefficient than the male series, indicating a greater inequality for males.
@@ 3.1 Cumulative distribution function transformation
    Univariate functional time series (UFTS) method: Applying eigendecomposition to $\bm{Z}^{s}=(\bm{Z}^{s}_{1},\dots,\bm{Z}^{s}_{n})^{\top}$ and $\bm{Z}^{s}_{1}=(Z^{s}_{1,1},\dots,Z^{s}_{n,111})$, it gives
@@
    where $\widehat{\eta}^{s}_{n+h|n,k}$ denotes the $h$-step-ahead univariate time-series forecast of the principal component scores. Among the univariate time-series methods, we consider the exponential smoothing method of Hyndman et al. (2008) for forecasting scores. In Figure 5, we present the 20 steps ahead forecasts of CDFs from 2023 to 2042 using the UFTS method.
    Figure 5: The 20 steps ahead forecasts of CDFs from 2023 to 2042 were obtained by the UFTS method. While the data from the distant past are shown in grey, the forecasts are shown in rainbow colours.
@@
    where $\Delta$ represents the first-order differencing. In Figure 6, we display the 20 steps ahead forecasts of life-table death counts.
    Figure 6: The 20 steps ahead forecasts of the life-table death counts from 2023 to 2042.
@@ 3.2 Centered log-ratio transformation
    Compute the simple mean function, which can be estimated by a simple average
@@ 4.1 Expanding window
An expanding window analysis of a time series model is commonly used to assess model and parameter stability over time, and prediction accuracy. The expanding window analysis determines the constancy of a model’s parameter by computing parameter estimates and their forecasts over an expanding window through the sample (for details Zivot & Wang, 2006, pp. 313-314). Using the first 32 observations from 1975 to 2006 in the Japanese female and male age-specific life-table death counts, we produce one- to 16-step-ahead forecasts. Through an expanding-window approach, we re-estimate the parameters in the time series forecasting models using the first 33 observations from 1975 to 2007. Forecasts from the estimated models are then produced for one- to 15-step-ahead point and interval forecasts. We iterate this process by increasing the sample size by one year until reaching the end of the data period in 2022. This process produces 16 one-step-ahead forecasts, 15 two-step-ahead forecasts, $\dots$, and one 16-step-ahead forecast. We compare these forecasts with the holdout samples to determine the accuracy of the out-of-sample forecast.
@@ 4.3 Interval forecast evaluation
$\text{ECP}_{\alpha}(h)=1-\frac{\sum^{16}_{\xi=h}\sum^{111}_{x=1}\left[\mathds{1}(d_{n+\xi,x}<\widehat{d}_{n+\xi|n,x})+\mathds{1}(d_{n+\xi,x}>\widehat{d}_{n+\xi|n,x}^{\text{ub}})\right]}{111\times(17-h)},$
@@ 4.4 Comparisons of point forecast accuracy
In Table LABEL:tab:1, we present the one- to 16-step-ahead point forecast errors of the life-table death counts in Japan. The errors are measured by the symmetrised Kullback-Leibler Divergence (KLD) and Jensen-Shannon Divergence (JSD), which are commonly used for evaluating density forecasts. Since the errors were minuscule, we multiplied them by 100 to highlight the differences among the methods. Between the CDF and clr transformations, the CDF transformation coupled with the MLFTS method produces the smallest errors for almost all forecast horizons. Between the two methods of selecting the number of components, using $K=6$ proves advantageous, confirming the early finding of Hyndman et al. (2013). Additionally, between the female and male series, the latter yields smaller errors.
Table 1: Comparison of the point forecast errors (×100)($\times$ 100)( × 100 ) of the age distribution of deaths between the CDF and clr transformations, when the number of components is fixed at $K=6$. Among the CDF transformation, we also consider the UFTS, MFTS and MLFTS methods.
@@ 5 Application to a single-premium temporary immediate annuity
Using the CDF transformation, we obtain forecasts of age-specific life-table death counts and then determine the corresponding age-specific survival probabilities. In Figure 7, we display the age-specific life-table death count forecasts from 2023 to 2072 for Japanese females and males.
@@
where $B(0,\tau$ is the $\tau$-year bond price and ${}_{\tau}p_{x}$ denotes the survival probability.
In Table 4, to provide an example of the annuity calculations, we compute the point estimate of the annuity prices for various ages and maturities for female and male policyholders in Japan. We assume a constant interest rate at $\eta=0.25\%$ (https://tradingeconomics.com/japan/interest-rate) and hence zero-coupon bond is given as $B(0,\tau)=\exp^{-\eta\tau}$. In Appendix D, we present the best estimate of annuity prices under a constant interest rate of $3\%$. Although the difference in annuity price might appear small, any mispricing can result in a significant risk when considering a large contract value and annuity portfolio.
Table 4: Estimates of annuity prices with different ages and maturities ($T$) for female and male policyholders residing in Japan. These estimates are based on forecast mortality from 2023 to 2072 and a constant interest rate $\eta=0.25\%$. We consider only contracts with maturity so that age + maturity $\leq 110$.
@@ 6 Conclusion
There are a few ways in which this paper could be extended, and we briefly discussed three. First, a robust MLFTS method may be considered in the presence of outlying years. Second, the methodology can be applied to calculate other annuity prices, such as the whole-life immediate or deferred annuity. Finally, we may also consider a cohort life table that depicts the life history of a specific group of individuals (see, e.g., Hatzopoulos & Haberman, 2015; Basellini et al., 2020), but is dependent on projected mortality rates for those cohorts born more recently.
@@ Appendix B Additional results of interval forecast accuracy comparison
In Table LABEL:tab:Appendix_2, we present the comparison of the interval forecast accuracy between the CDF and clr transformation at the 80% nominal coverage probability when the number of components is selected by the EVR criterion. For the female data, the clr transformation produces the smallest averaged interval score, while the CDF transformation coupled with the MLFTS method gives the smallest averaged CPD. For the male data, the clr transformation gives the smallest interval forecast errors.
@@ Appendix D Annuity price calculation under an interest rate $\eta=3\%$
In Table 9, we display the annuity price calculation for female and male policyholders residing in Japan under a constant interest rate of $\eta=3\%$ in line with the US, UK and Australia. In contrast to $\eta=0.25\%$, the following annuity premiums are lower for each age and maturity period.

### v2

@@ 1 Introduction
In demography, the age distribution of deaths provides important insights into longevity and lifespan variability that cannot be grasped directly from either the central mortality rate or the survival function. For modelling the age distribution of deaths, Oeppen (2008) demonstrated that using compositional data analysis (CoDa) to forecast the age distribution of death counts does not lead to more pessimistic results than forecasting age-specific mortality. The pessimistic mortality rate forecasts may also be due to the linear extrapolation (constant trend) in the well-known Lee-Carter model; in that case, the mortality rate forecasts are higher than the actual ones. Thus, the age distribution of death counts is more suitable than the central mortality rates for computing period life expectancy and annuity values. In addition to providing an informative description of the mortality experience of a population, the age-at-death distribution yields readily available information on the “central longevity indicators” (e.g., mean, median and modal age at death, see Cheung et al. (2005) and Canudas-Romo (2010)) and lifespan variability (e.g., Robine, 2001; Vaupel et al., 2011; van Raalte & Caswell, 2013; van Raalte et al., 2014; Aburto & van Raalte, 2018).
@@
Using the Japanese age- and sex-specific life-table death counts from 1975 to 2022, we evaluate and compare one- to 16-step-ahead point and interval forecast accuracy between the CoDa and CDF transformations. To evaluate point forecast accuracy, we use the Kullback-Leibler divergence and the Jensen-Shannon divergence. To assess the interval forecast accuracy, we utilise the interval score of Gneiting & Raftery (2007) and Gneiting & Katzfuss (2014); see Section 4 for details. The improved forecast accuracy of life-table death counts is of great importance to demographers and actuaries for determining remaining age-specific life expectancies and to actuaries for pricing temporary annuities for various ages and maturities.
@@ 2 Age- and sex-specific life-table death counts in Japan
In many developed nations, such as Japan, increases in longevity risk and an aging population has led to concerns regarding the sustainability of the pension, health and aged care systems (see, e.g., Coulmas, 2007; OECD, 2020). Japan has one of the highest average life expectancies in the world, with exceptional longevity in Okinawa (Poulain, 2011). Japan is also one of the countries that have freely available subnational mortality data, alongside Australia, Canada, France and the US.
In Figure 1, we display Japanese age- and sex-specific life-table death counts from 1975 to 2022, obtained from the Japanese Mortality Database (2024). We use data from the period after the First and Second World Wars to obtain a more stable parameter estimate from the historical data. We study life-table death counts, where the life-table radix (i.e., a population experiencing 100,000 births annually) is fixed at 100,000 at age 0 while the remaining number of people alive is 0 in the last age group for each year. For the life-table death counts, there are 111 ages, and these are ages 0, 1, …, 109, 110+. Due to rounding, there are potentially zero counts for age 110+ at some years. To overcome this problem, we work with the probability of dying (i.e., $q_{x}$) and the life-table radix to recalculate our estimated death counts (up to six decimal places). In doing so, we obtain more precise death counts than the ones reported in the Japanese Mortality Database (2024).
@@
Figure 1: Rainbow plots of the age distribution of deaths from 1975 to 2022 in a single-year group. The life-table radix is 100,000 for each year. The life-table death counts in the oldest years are shown in red, while the most recent years are in violet. Curves are ordered chronologically according to the rainbow colours.
@@
Lifespan variability can be assessed using measures like the interquartile range or the Gini coefficient (see Wilmoth & Horiuchi, 1999; van Raalte & Caswell, 2013; Debón et al., 2017). In life-table death counts, a Gini coefficient of 0 represents perfect inequality among ages, while 1 indicates perfect equality, with deaths occurring at the same age. Figure 2 shows higher Gini coefficients for females, reflecting greater inequality among males. While not our focus, quantifying and examining temporal changes in distribution spread is a key advantage of modeling age distributions of deaths.
@@ 3.1 Cumulative distribution function transformation
    Univariate functional time series (UFTS) method: With a set of unconstrained data $\bm{Z}^{s}=(\bm{Z}^{s}_{1},\dots,\bm{Z}^{s}_{n})^{\top}$ and $\bm{Z}^{s}_{1}=(Z^{s}_{1,1},\dots,Z^{s}_{n,111})$, we consider a functional time series forecasting technique (see, e.g., Kokoszka & Reimherr, 2017; Haghbin & Maadooliat, 2024). It is assumed that the observations $\bm{Z}^{s}$ are elements of the Hilbert space equipped with an inner product. Each function is a square-integrable function with a finite variance.
    Assuming the existence of the mean and finite variance, its non-negative definite covariance function is given by
    $\text{Cov}^{s}(x,y)$ $:=\text{Cov}(\bm{Z}_{x}^{s},\bm{Z}_{y}^{s})$
    $=E(Z_{x}^{s}-\mu^{s}_{x})(Z_{y}^{s}-\mu^{s}_{y}),$
    where $\mu^{s}_{x}=\frac{1}{n}\sum^{n}_{t=1}Z_{t,x}^{s}$, $x$ or $y$ represents a continuum in a function support range $\mathcal{I}\in R$ and $R$ is the real line. Via Mercer’s lemma, there exists a set of orthonormal sequence $\psi^{s}_{k,x}$ of continuous function and a non-increasing sequence $\lambda_{k}$ of positive numbers, such that
    $\text{Cov}^{s}(x,y)=\sum^{\infty}_{\lambda=1}\lambda_{k}^{s}\psi_{k,x}^{s}\psi_{k,y}^{s},$
    where $\lambda_{k}^{s}$ and $\psi_{k,x}^{s}$ denote the $k$th eigenvalue and eigenfunction, respectively. The Karhunen-Loève expansion of a stochastic process $Z^{s}_{t,x}$ can be expressed as
@@
    where $\widehat{\eta}^{s}_{n+h|n,k}$ denotes the $h$-step-ahead univariate time-series forecast of the principal component scores. Among the univariate time-series methods, we consider the exponential smoothing method of Hyndman et al. (2008) for forecasting scores. In Figure 5, we present the 16 steps ahead forecasts of CDFs from 2023 to 2038 using the UFTS method.
    Figure 5: The 16 steps ahead forecasts of CDFs from 2023 to 2038 were obtained by the UFTS method. While the data from the distant past are shown in grey, the forecasts are shown in rainbow colours.
@@
    where $\Delta$ represents the first-order differencing. In Figure 6, we display the 16 steps ahead forecasts of life-table death counts.
    Figure 6: The 16 steps ahead forecasts of the life-table death counts from 2023 to 2038.
@@ 3.2 Centered log-ratio transformation
    Compute the geometric mean function, which can be estimated by a simple average
@@ 4.1 Expanding window
An expanding window analysis of a time series model is commonly used to assess model and parameter stability over time, and prediction accuracy. The expanding window analysis determines the constancy of a model’s parameter by computing parameter estimates and their forecasts over an expanding window through the sample (for details Zivot & Wang, 2006, pp. 313-314). Using the first 32 observations from 1975 to 2006 in the Japanese female and male age-specific life-table death counts, we produce one- to 16-step-ahead forecasts. Through an expanding-window approach, we estimate the parameters in the time series forecasting models using the first 33 observations from 1975 to 2007. Forecasts from the estimated models are then produced for one- to 15-step-ahead point and interval forecasts. We iterate this process by increasing the sample size by one year until reaching the end of the data period in 2022. This process produces 16 one-step-ahead forecasts, 15 two-step-ahead forecasts, $\dots$, and one 16-step-ahead forecast. We compare these forecasts with the holdout samples to determine the accuracy of the out-of-sample forecast.
@@ 4.3 Interval forecast evaluation
$\text{ECP}_{\alpha}(h)=1-\frac{\sum^{16}_{\xi=h}\sum^{111}_{x=1}\left[\mathds{1}(d_{n+\xi,x}<\widehat{d}^{\text{lb}}_{n+\xi|n,x})+\mathds{1}(d_{n+\xi,x}>\widehat{d}_{n+\xi|n,x}^{\text{ub}})\right]}{111\times(17-h)},$
@@ 4.4 Comparisons of point forecast accuracy
In Table LABEL:tab:1, we present the one- to 16-step-ahead point forecast errors of the life-table death counts in Japan. The errors are measured by the symmetrised Kullback-Leibler Divergence (KLD) and Jensen-Shannon Divergence (JSD), which are commonly used for evaluating density forecasts. Since the errors were minuscule, we multiplied them by 100 to highlight the differences among the methods. Between the CDF and clr transformations, the CDF transformation coupled with the MLFTS method produces the smallest errors for almost all forecast horizons. Between the two methods of selecting the number of components, using $K=6$ proves advantageous, confirming the early finding of Hyndman et al. (2013). Additionally, between the female and male series, the latter yields smaller errors. For male data, the clr transformation performs better at longer horizons. Although that could be due to data randomness, the finding motivates us to consider a model-averaging approach by assigning weights to forecasts obtained from the two transformations.
Table 1: Comparison of the point forecast errors (×100)($\times$ 100)( × 100 ) of the age distribution of deaths between the CDF and clr transformations, when the number of components is fixed at $K=6$. Among the CDF transformation, we also consider the UFTS, MFTS and MLFTS methods. The smallest errors are shown in bold.
@@ 5 Application to a single-premium temporary immediate annuity
Using the CDF transformation, we apply the multilevel functional time series method to obtain forecasts of age-specific life-table death counts and then determine the corresponding age-specific survival probabilities. In Figure 7, we display the age-specific life-table death count forecasts from 2023 to 2072 for Japanese females and males.
@@
where $B(0,\tau)$ is the $\tau$-year bond price and ${}_{\tau}p_{x}$ denotes the survival probability.
Table 4: Estimates of annuity prices with different ages and maturities ($T$) for female and male populations residing in Japan. These estimates are based on forecast mortality from 2023 to 2072 and a constant interest rate $\eta=0.25\%$. We consider only contracts with maturity so that age + maturity $\leq 110$.
@@
In Table 4, to provide an example of the annuity calculations, we compute the point estimate of the annuity prices for various ages and maturities for female and male populations in Japan. We assume a constant current interest rate at $\eta=0.25\%$ (https://tradingeconomics.com/japan/interest-rate) and hence zero-coupon bond is given as $B(0,\tau)=\exp^{-\eta\tau}$. In Appendix D, we present the best estimate of annuity prices under a constant interest rate of $3\%$. Although the difference in annuity price might appear small, any mispricing can result in a significant risk when considering a large contract value and annuity portfolio.
@@ 6 Conclusion
There are a few ways in which this paper could be extended, and we briefly discussed four.
  1. 1)
    A robust MLFTS method may be considered in the presence of outlying years.
  2. 2)
    A model-averaging approach can improve accuracy for forecasting the age distribution of death counts.
  3. 3)
    The methodology can be applied to calculate other annuity prices, such as the whole-life immediate or deferred annuity.
  4. 4)
    We may also consider a cohort life table that depicts the life history of a specific group of individuals (see, e.g., Hatzopoulos & Haberman, 2015; Basellini et al., 2020), but is dependent on projected mortality rates for those cohorts born more recently.
@@ Appendix B Additional results of interval forecast accuracy comparison
In Table LABEL:tab:Appendix_2, we present the comparison of the interval forecast accuracy between the CDF and clr transformation at the 80% nominal coverage probability when the number of components is selected by the EVR criterion. For the female data, the clr transformation produces the smallest averaged interval score, while the CDF transformation coupled with the MLFTS method gives the smallest averaged CPD. For the male data, the clr transformation gives the smallest interval forecast errors. The mixed results motivate us to consider a model-averaging approach by assigning weights to forecasts obtained from the two transformations.
@@ Appendix D Annuity price calculation under an interest rate $\eta=3\%$
In Table 9, we display the annuity price calculation for female and male populations residing in Japan under a constant interest rate of $\eta=3\%$ in line with the US, UK and Australia. In contrast to $\eta=0.25\%$, the following annuity premiums are lower for each age and maturity period.